OpenStudy (helpblahblahblah):

Medal+ fan plz help :) Which shows the expressions in the order they would appear on a number line from least to greatest? A. 17 over 9, square root of 6, square root of 15, square root of 30, 3 to the power of 3 B. 3 to the power of 3, square root of 30, square root of 15, square root of 6, 17 over 9 C. 3 to the power of 3, square root of 6, square root of 15, 17 over 9, square root of 30 D. 17 over 9, square root of 30, square root of 15, 3 to the power of 3, square root of 6

OpenStudy (helpblahblahblah):

@Opcode

OpenStudy (helpblahblahblah):

@nightwielder

OpenStudy (helpblahblahblah):

@esamalaa

OpenStudy (mathmath333):

\(\large\tt \color{black}{\dfrac{17}{9},\sqrt6,\sqrt{15}, \sqrt{30}, 3^3}\) \(\large\color{red}{square~~ all ~~of~~ them}\) \(\large\tt \color{black}{(\dfrac{17}{9})^2,6,15, 30, (3^3)^2}\) \(\large\tt \color{black}{3.56<6<15< 30< (27)^2}\)

OpenStudy (mathmath333):

so option a is correct i think

OpenStudy (helpblahblahblah):

thxxxxx :D